OpenStudy (anonymous):

f(x)=3x^2-3x, f (0)=, f (-1)=, f (2)=

OpenStudy (anonymous):

plug, then chug.

OpenStudy (anonymous):

what you mean plug than chug

OpenStudy (anonymous):

Replace x with 0 and evaluate the function. Then do the same thing for -1 and 2. If you are sufficiently skilled in math to be evaluating powers (\(x^2\)) you should be sufficiently skilled in substituting values for a variable. Did you have a specific question on how to arrive at an answer here?

OpenStudy (anonymous):

Yes I have never done algebra this is my first year in algebra thats y I said how do you plug and chug

OpenStudy (anonymous):

f(0) means take 0 and replace x with it. Then evaluate the function to find the result.

OpenStudy (anonymous):

\[3(0)^2-3(0) = 3*0 - 3*0 = 0-0 = 0\]

OpenStudy (anonymous):

Now do the same thing for -1 and 2.

OpenStudy (anonymous):

ok so I use the 0 with the -1 and 2

OpenStudy (anonymous):

No. You use the function with -1 and 2.

OpenStudy (anonymous):

I am so sry I am still confused, I guess I just skip over it

OpenStudy (anonymous):

No, don't skip over it. Figure out what is confusing to you. Is English your native language, cause it's not clear to me which part of my explanation doesn't make sense. \[f(x) = 3x^2-3x\] Now if I told you that x = 0 what would you do to find f(x) ?

OpenStudy (anonymous):

will the answer be 0,3,6

OpenStudy (anonymous):

Close.. Should be 0, 6, and 6.

OpenStudy (anonymous):

Yea i didn't do that f(1) right

OpenStudy (anonymous):

it's f(-1)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

At least according to what you wrote initially.

OpenStudy (anonymous):

Thanks so much I think I got it can I be a fan

OpenStudy (anonymous):

So.. \[f(-1) = 3(-1)^2 - 3(-1)\]

OpenStudy (anonymous):

Of course!

OpenStudy (anonymous):

Thanks so much